If a man buys two products by the pound and pays $0.26 for one and $0.19 for the other. His total amount payed is $10,530.03 and his total Weight is 42,110 pounds. What is the weight of product bought at $0.26 and what is the weight of product bought at $0.19 ??? Please help!
Found 2 solutions by stanbon, checkley77:
Answer by stanbon(48568)

If a man buys two products by the pound and pays $0.26 for one and $0.19 for the other. His total amount payed is $10,530.03 and his total Weight is 42,110 pounds. What is the weight of product bought at $0.26 and what is the weight of product bought at $0.19 ???
------------------------------
Weight Equation: x + y = 42110
Value equation: 0.26x + 0.19y = 10530.03
------------------------------
Multiply thru the Value Eq. by 100
Multiply thru the Weight Eq. by 26
---
26x + 19y = 1053003
26x + 26y = 26*42110
---------------------
Subtract the 1st from the 2nd and solve for "y":
7y = 41857
y = 5979.5714 lbs (Weight of product bought at $0.19)
==========================================================
Cheers,
Stan H.

Answer by checkley77(12569)

.26x+.19(42,110-x)=10,530.03
.26x+8,000.9-.19x=10,530.03
.07x=10,530.03-8,000.9
07x=2,529.13
x=2,529.13/.07
x=36,130.43 pounds of $.26 product.
42,110-36,130.43=5,979.57 pounds of $.19 product.
Proof:
.26*36,130.43+.19*5.979.57=10,530.03
9,393.91+1.136.12=10,530.03
10,530.03=10,530.03